OpenStudy (anonymous):

In 2005, four year old Jaiden planted a tree that was 3 feet tall.In 2010, the tree was 13 feet tall. Assuming the growth of the tree is linear, what was the rate of growth of the tree? Write an equation.

jigglypuff314 (jigglypuff314):

Hello @things1234567 and Welcome to OpenStudy! :) from this we can get points like (2005, 3) and (2010, 13) out of it rate = slope and the equation for slope is \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given points (x1, y1) and (x2, y2) knowing this much, we can plug in and get the rate that they want \[\frac{ 13-3 }{ 2010-2005 } =~~?\]can you solve that? :)